solve


dy/dx+4xy=8x

(6xy+2y2-5)dx+(3x2+4xy-6)dy=0 y(1)=2


1
Expert's answer
2021-07-28T15:48:57-0400

(a) Solve the separable equation


"\frac{d y(x)}{d x}+4 x y(x)=8x :"


Solve for "\frac{d y(x)}{d x} :" :

"\frac{d y(x)}{d x}=-4(-2 x+x y(x))"

Simplify:


"\frac{d y(x)}{d x}=x(-4 y(x)+8)"


Divide both sides by "-4 y(x)+8:"


"\frac{\frac{d y(x)}{d x}}{-4 y(x)+8}=x"


Integrate both sides with respect to "\boldsymbol{x}:"


"\int \frac{\frac{d y(x)}{d x}}{-4 y(x)+8} d x= \int x d x"


Evaluating this integrals:


"-\frac{1}{4} \log (-4 y(x)+8)=\frac{x^{2}}{2}+c_{1}"

where "c_{1}" is an arbitrary constant


Solve for "y(x):"


"y(x)=-\frac{1}{4} e^{-2\left(x^{2}+2 c_{1}\right)}+2"

By simplifying the arbitrary constants, we have:


"y(x)=-\frac{1}{4} e^{-2\left(x^{2}+c_{1}\right)}+2"

(b)


"(6xy+2y^2-5)dx+(3x^2+4xy-6)dy=0 \qquad y(1)=2"

The given differential equation is of the form:


"Mdx + Ndy =0"

With:


"M = 6xy+2y^2-5 \text{ and } N= 3x^2+4xy-6"

We proceed to check the exactness of the two equations. For exactness,


"\frac{\partial M}{\partial y}=\frac{\partial N}{\partial x}"

Thus:


"\frac{\partial M}{\partial y}=6x+4y=\frac{\partial N}{\partial x}"

We then proceed to solve the equation.


"\Phi_x = \int(6xy+2y^2-5)dx = 3x^2y+2xy^2-5x\\
\Phi_y= \int (3x^2+4xy-6)dy= 3x^2y+2xy^2-6y"

The general solution of the DE is:


"3x^2y+2xy^2-5x-6y=c"

Evaluating the general solution using the IVP "y(1)=2" :


"3(1)^2(2)+2(1)(2)^2-5(1)-6(2)=c\\
6+8-5-12=c\\
\therefore c=-3"

Thus the solution is:


"3x^2y+2xy^2-5x-6y=-3"